OpenStudy (anonymous):

4y-2 over 5 - y-4 over 5 = -3

OpenStudy (anonymous):

\[\frac{ 4y -2 }{ 5 } - \frac{ y -4 }{ 5 } = -3\]

OpenStudy (anonymous):

\[\frac{ 4y - 2-y -4 }{ 5 } = -3\]

OpenStudy (anonymous):

\[3y-6 = -15\]

OpenStudy (anonymous):

\[3y = -15 +6\]

OpenStudy (anonymous):

\[3y = -9\] \[y=-3\]

OpenStudy (anonymous):

ugh thank you!!!!